Z=a(x+log y)-x^2/2-bx

1
Expert's answer
2021-03-08T18:28:07-0500

"z(x,y)=a(x+\log y)-\frac{x^2}{2}-bx\\
z= ax +a\log y-\frac{x^2}{2}-bx"

for z(x,y):


"dz = \dfrac{\partial z}{\partial x}dx+\dfrac{\partial z}{\partial y}dy\\"

then;


"\dfrac{\partial z}{\partial x} = a - \frac{x^3}{6}-b\; \text{;} \qquad \dfrac{\partial z}{\partial y} = \frac{a}{y}"

Let:


"p = \dfrac{\partial z}{\partial x} \qquad \text{and } \qquad q=\dfrac{\partial z}{\partial y}"

then


"q = \frac{a}{y} \implies a =qy \quad \cdots (i)\\
p = a -\frac{x^3}{6}-b \qquad \cdots(ii)"

put (i) in (ii)


"p = qy -\frac{x^3}{6}-b\\
6p = 6qy - x^3 - 6b\\
6p-6qy+x^3+6b=0\\
6(p-qy)+x^3+6b=0\\
6(p-qy+b)+x^3=0\\"

The required PDE is therefore:


"6(p-qy+b)+x^3=0\\"